OpenStudy (anonymous):

A triangle has two side lengths of 8.1 and 1.4 if the length of the third side is a even integer its length must be A) 2 B) 4 C) 6 D) 8 E) 10

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

I'll assume 8.1 is the hyp. 8.1^2 = 1.4^2 + ?^2 65.61 = 1.96 + ?^2 ?^2 = 63.65 ? = 8